Suppose that an investment promises to pay a nominal 9.6 percent annual rate of interest. What is the effective annual interest rate on this investment assuming that interest is compounded (a) annually? (b) semiannually? (c) quarterly? (d) monthly? (e) daily (365 days)? (f ) continuously? (Note: Report your answers accurate to four decimal places – e.g., 0.0987 or 9.87%.)


1
Expert's answer
2021-05-20T14:44:54-0400

Active yearly interest rate = (1 + I/n)n - 1

i = Nominal interest rate

n = Quantity of stages.

a) Per annum.

= (1+9.6%/1)1 - 1

=0.096

b) Semi-annually

(1+9.6%/2)2 - 1

= 0.0983

c) Quarterly

(1+9.6%/4)4 - 1

= 0.0995

d) monthly

(1+9.6%/12)12 - 1

= 0.1003

e) Day-to-day

(1+9.6%/365)365 - 1

= 0.1007

f) Incessant is computed by rising the number "e" (2.71828) to the command of the interest rate and deducting by one.

=2.71828^0.096-1

= 0.1008

22). Multiple Yearly Development Rate

CAGR = [(commencement balance or termination balance)1/n - 1]

[(2000/1000)1/14 - 1]

= 5.08%

[(4000/2000)1/8 - 1]

= 9.05%

[(8000/4000)1/2 - 1]

= 41.42%